OpenStudy (anonymous):

2a^2 - 13a + 15 / 8a^2 -12a * 6a - 4a^2 / a^2 -10a + 25

jhonyy9 (jhonyy9):

(2a-3)(a-5)/4a(2a-3) * (-2a(2a-3))/((a-5)*(a-5))= =(a-5)/4a * (-2a(2a-3))/((a-5)*(a-5))= -(2a-3)/(2(a-5))=(3-2a)/(2a-10)